### v2

@@ Appendix D Convergence proofs for the inexact PPM in Section 5
The optimality condition of 20 and criteria $\text{A}^{\prime}$ implies that there exist $v_{k}\in\partial f(x_{k+1})$ and $\theta_{k}\in\mathbb{R}^{n}$ such that
@@
This improvement holds for all $k\geq 0$. Summing above inequality from 00 to $k-1$ yields
@@
$2\sum_{j=0}^{k-1}c_{j}\min_{i=0,\ldots,k}f(x_{i})-f(x^{\star})$ $\leq 2\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f(x^{\star}))$
$\leq\|x_{0}-x^{\star}\|^{2}+2D\sum_{j=0}^{k-1}\epsilon_{j}.$
@@
Remark D.1 (Average/weighted iterate).
@@
$f(\bar{x}_{k})-f^{\star}\leq\frac{\sum_{j=0}^{k-1}(f(x_{j+1})-f^{\star})}{k}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ f(\tilde{x}_{k})-f^{\star}\leq\frac{\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f^{\star})}{\sum_{j=0}^{k-1}c_{j}}.$
The proof in Theorem 5.2 borrows some proof idea from [47, Theorem 3] with the fact that $\text{A}^{\prime}$ implies the existence of $\theta_{k}\in\mathbb{R}^{n}$ such that $\mathrm{dist}(0,H_{k}(x_{k+1}))={\|\theta_{k}\|}/{c_{k}}\leq\epsilon_{k}$. ∎
@@
The proof is an adaptation of [33, Section 2]. For completeness, we reprove it here. Fix a $x\in\mathbb{R}^{n}$. From the update 12, we know
@@
$\forall x\in\mathbb{R}^{n},\leavevmode\nobreak\ \|x-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ $\leq\|x-\Pi_{S}(x)\|+\|\Pi_{S}(x)-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
@@
$\leavevmode\nobreak\ \|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\leavevmode\nobreak\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|+\|\text{prox}_{c_{k}f}(x_{k})-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\leavevmode\nobreak\ \delta_{k}\|x_{k+1}-x_{k}\|+\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S)$
$\leq$ $\leavevmode\nobreak\ \delta_{k}\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|+\delta_{k}\|x_{k}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
@@
Finally, using the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ and 38 for $x=x_{k}$ completes the proof.

### v3

@@ Appendix D Convergence proofs for the inexact PPM in Section 5
The optimality condition of 23 and criteria $\text{A}^{\prime}$ implies that there exist $v_{k}\in\partial f(x_{k+1})$ and $\theta_{k}\in\mathbb{R}^{n}$ such that
@@
This improvement holds for all $k\geq 0$. Summing above inequality from 0 to k1k-1 yields
@@
$2\sum_{j=0}^{k-1}c_{j}\min_{i=0,\ldots,k}f(x_{i})-f(x^{\star})$ $\leq 2\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f(x^{\star}))\leq\|x_{0}-x^{\star}\|^{2}+2D\sum_{j=0}^{k-1}\epsilon_{j}.$
@@
Remark 3 (Average/weighted iterate).
@@
$f(\bar{x}_{k})-f^{\star}\penalty 10000\ \text{and}\penalty 10000\ f(\tilde{x}_{k})-f^{\star}\leq\frac{\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f^{\star})}{\sum_{j=0}^{k-1}c_{j}}.$
The proof in Theorem 5.2 borrows some proof idea from [lu2023unified, Theorem 3] with the fact that $\text{A}^{\prime}$ implies the existence of $\theta_{k}\in\mathbb{R}^{n}$ such that $\mathrm{dist}(0,H_{k}(x_{k+1}))=\frac{\|\theta_{k}\|}{c_{k}}\leq\epsilon_{k}$. ∎
@@
The proof is an adaptation of [luque1984asymptotic, Section 2]. For completeness, we reprove it here. Fix a $x\in\mathbb{R}^{n}$. From the update 15, we know
@@
$\forall x\in\mathbb{R}^{n},\penalty 10000\ \|x-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ $\leq\|x-\Pi_{S}(x)\|+\|\Pi_{S}(x)-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
@@
$\penalty 10000\ \|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\penalty 10000\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|+\|\text{prox}_{c_{k}f}(x_{k})-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\penalty 10000\ \delta_{k}\|x_{k+1}-x_{k}\|+\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S)$
$\leq$ $\penalty 10000\ \delta_{k}\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|+\delta_{k}\|x_{k}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
@@
Finally, using the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ and 41 for $x=x_{k}$ completes the proof.
Declarations
[…]